# Limited-Range Multilinear Off-Diagonal Extrapolation and Weighted Transference Principle

### v1

@@ 3 Limited-Range Multilinear Off-Diagonal Mixed-Norm Extrapolation
$\|\widetilde{T}\vec{f}\|_{L^{\tilde{\mathbf{q}}}_{\tilde{\boldsymbol{{v}}}}(\Omega_{\Lambda_{2}})}\leq\kappa^{(\ell-1)/\alpha}\tilde{\phi}((\tilde{\phi}_{j}(\kappa^{\prime},[\tilde{\boldsymbol{{w}}}_{j},\tilde{\boldsymbol{{v}}}_{j}]_{\tilde{\mathbf{s}}_{j},\tilde{\mathbf{r}}_{j}}))_{j\in J})\prod_{j\in J}\|\widetilde{S}_{j}f_{j}\|_{L^{\tilde{\mathbf{p}}_{j}}_{\tilde{\boldsymbol{{w}}}_{j}}(\Omega_{\Lambda_{2}})},$ (20)
@@
    The proof of Theorem 3.4 allows to track the exact definition of $\phi_{\vec{s_{0}},\vec{r_{0}},\vec{\boldsymbol{\gamma}}}$ recursively. In particular, assume that there is $c:(1,\infty]\to[1,\infty)$ such that for all $p\in(1,\infty]$, $i\in I$, and weights w{w} on $\Omega_{\Lambda_{i}}$ with ${w}]_{p}^{{\mathcal{U}}_{\Lambda_{i}}}<\infty$, we have that ${\mathcal{M}}_{{\mathcal{U}}_{\Lambda_{i}}}$ is bounded from $L^{p}_{{w}}(\Omega_{\Lambda_{i}})$ to $L^{p}_{{v}}(\Omega_{\Lambda_{i}})$ with
@@
    Then if $\vec{\mathbf{s}},\vec{\mathbf{r}}\in(0,\infty)^{\ell\times m}$, w=v{w}={v}, and $\kappa\in(1,\infty)$, an iteration of the bound (11) in Remark 2.7 shows that we may choose $\phi_{\vec{s_{0}},\vec{r_{0}},\vec{\boldsymbol{\gamma}}}([\vec{\boldsymbol{{w}}},\vec{\boldsymbol{{v}}}]_{\vec{\mathbf{s}},\vec{\mathbf{r}}}):=\kappa^{\ell/\alpha}\phi(\vec{C})$ with
@@
    Theorem 3.4 also contains a weak-type extrapolation result by an argument of Grafakos and Martell [8]: One may exchange $\|T\vec{f}\|_{L^{q_{0}}_{{w}_{0}}(\Omega)}$ by $\|T\vec{f}\|_{L^{q_{0},\infty}_{{w}_{0}}(\Omega)}$ in (16) if one also exchanges $\|T\vec{f}\|_{L^{\mathbf{q}}_{\boldsymbol{{w}}}(\Omega)}$ by $\|T\vec{f}\|_{L^{\mathbf{q},\infty}_{\boldsymbol{{w}}}(\Omega)}$ in (17). Here, the weak mixed-norm space $L^{\mathbf{p},\infty}(\Omega)$ is given as all $f\in L^{0}(\Omega)$ such that
@@
    where $E_{\lambda}:=\{x\in\Omega\ \lvert\ |f|>\lambda\}$, and $f\in L^{\mathbf{p},\infty}_{\boldsymbol{{w}}}(\Omega)$ if $f{w}\in L^{\mathbf{p},\infty}(\Omega)$, where w{w} is the tensor weight induced by $\boldsymbol{{w}}$. To see the weak-type extrapolation, consider for $\lambda>0$ the set $E_{\lambda}:=\{x\in\Omega\ \lvert\ |T\vec{f}(x)|>\lambda\}$ and define $T_{\lambda}f:=\lambda\mathds{1}_{E_{\lambda}}$. Applying Theorem 3.4 with $T\mapsto T_{\lambda}$, the assertion is obtained in the same way as in [8, Theorem 6.1].
@@
$\mathcal{T}\subseteq\{T:\prod_{j\in J}\bigcup_{({w}_{0j},{v}_{0j})\in W_{j}}S_{j}^{-1}(L^{p_{0j}}_{{w}_{0j}}(\Omega))\to L^{0}(\Omega)\}.$
If there is a function $\phi:\mathbb{R}^{m}\to\mathbb{R}$ increasing in each component such that for all $T\in{\mathcal{T}}$, $(\vec{{w}}_{0},\vec{{v}}_{0})\in\vec{W}:=W_{1}\times\cdots\times W_{m}$, and $\vec{f}\in\prod_{j\in J}S_{j}^{-1}(L^{\vec{p}_{0}}_{\vec{{w}}_{0}}(\Omega))$ it holds
$\|T\vec{f}\|_{L^{q_{0}}_{{v}_{0}}(\Omega)}\leq\phi([\vec{{w}}_{0},\vec{{v}}_{0}]_{(\vec{s}_{0},\vec{r}_{0})})\prod_{j\in J}\|S_{j}f_{j}\|_{L^{p_{0j}}_{{w}_{0j}}(\Omega)},$ (21)
then for all $T\in\mathcal{T}$ and all $\vec{f}\in\prod_{j\in J}S_{j}^{-1}(L^{\mathbf{p}_{j}}_{\boldsymbol{{w}}_{j}}(\Omega))$, $T\vec{f}$ is well defined, and
$\|(T_{k}(f_{1k},\ldots,f_{jk}))_{k\in\mathbb{N}}\|_{L^{\mathbf{q}}_{\boldsymbol{{v}}}(\Omega;\ell^{u_{0}})}\leq\phi_{\vec{s_{0}},\vec{r_{0}},\vec{\boldsymbol{\gamma}}}([\vec{\boldsymbol{{w}}},\vec{\boldsymbol{{v}}}]_{\vec{\mathbf{s}},\vec{\mathbf{r}}})\prod_{j\in J}\|(f_{jk})_{k\in\mathbb{N}}\|_{L^{\mathbf{p}_{j}}_{\boldsymbol{{w}}_{j}}(\Omega;\ell^{u_{0j}})},$
for all $(T_{k})_{k\in\mathbb{N}}\subseteq\mathcal{T}$ and $(f_{jk})_{k\in\mathbb{N}}\subseteq V_{j}$ with $(S_{j}f_{jk})_{k\in\mathbb{N}}\in L^{\mathbf{p}_{j}}_{\boldsymbol{{w}}_{j}}(\Omega;\ell^{u_{0j}})$ for all $j\in J$. The function $\phi_{\vec{s_{0}},\vec{r_{0}},\vec{\boldsymbol{\gamma}}}$ can be taken identical to the one in Theorem 3.4.
@@
$\widetilde{V}_{j}:=\{(f_{jk})_{k\in\mathbb{N}}\subseteq V_{j}\ \lvert\ (S_{j}f_{jk})_{k\in\mathbb{N}}\in\bigcup_{({w}_{0j},{v}_{0j})\in W_{j}}L^{p_{0j}}_{{w}_{0j}}(\Omega;\ell^{u_{0j}})\}$
@@
We claim that $\widetilde{S}_{j}^{-1}(L^{\mathbf{p}_{j}}_{\boldsymbol{{w}}_{j}}(\Omega))=\{(f_{jk})_{k\in\mathbb{N}}\subseteq V_{j}\ \lvert\ (S_{j}f_{jk})_{k\in\mathbb{N}}\in L^{\mathbf{p}_{j}}_{\boldsymbol{{w}}_{j}}(\Omega;\ell^{u_{0j}})\}$ for all $j\in J$. The inclusion $\subseteq$” follows directly from the definition of $\widetilde{S}_{j}$. For $\supseteq$” let $(f_{jk})_{k\in\mathbb{N}}$ be in the set on the right-hand side and consider the function Then $h:=\|(S_{j}f_{jk})_{k\in\mathbb{N}}\|_{\ell^{u_{0j}}}\in L^{\mathbf{p}_{j}}_{\boldsymbol{{w}}_{j}}(\Omega)$. Then $h\in\bigcup_{({w}_{0j},{v}_{0j})\in W_{j}}L^{p_{0j}}_{{w}_{0j}}(\Omega)$ by Lemma 3.3, that is $(f_{jk})_{k\in\mathbb{N}}\subseteq\widetilde{V}_{j}$. Hence we may apply $\widetilde{S}_{j}$ to $(f_{jk})_{k\in\mathbb{N}}$ to see that $\widetilde{S}_{j}(f_{jk})_{k\in\mathbb{N}}=h\in L^{\mathbf{p}_{j}}_{\boldsymbol{{w}}_{j}}(\Omega)$, i.e. $(f_{jk})_{k\in\mathbb{N}}\in\widetilde{S}_{j}^{-1}(L^{\mathbf{p}_{j}}_{\boldsymbol{{w}}_{j}}(\Omega))$.
@@
$\|(S_{j}f_{jk})_{k}\|_{\ell^{u_{0j}}(L^{p_{0j}}_{{w}_{0j}}(\Omega))}\leq\|\widetilde{S}_{j}f_{j}\|_{L^{p_{0j}}_{{w}_{0j}}(\Omega)}.$
@@
$\|\widetilde{T}\vec{f}$ $\|_{L^{q_{0}}_{\vec{{w}}_{0}}(\Omega)}\leq\|(T_{k}\vec{f}_{k})_{k}\|_{\ell^{u_{0}}(L^{q_{0}}_{\vec{{w}}_{0}}(\Omega))}\leq\phi([{\vec{{w}}_{0}}]_{(\vec{s}_{0},\vec{r}_{0})})\left\|\prod_{j=1}^{m}\|(S_{j}f_{jk})_{k}\|_{L^{p_{0j}}_{{w}_{0j}}(\Omega)}\right\|_{\ell^{u_{0}}}$
$\leq\phi([{\vec{{w}}_{0}}]_{(\vec{s}_{0},\vec{r}_{0})})\prod_{j=1}^{m}\|(S_{j}f_{jk})_{k}\|_{\ell^{u_{0j}}(L^{p_{0j}}_{{w}_{0j}}(\Omega))}\leq\phi([{\vec{{w}}_{0}}]_{(\vec{s}_{0},\vec{r}_{0})})\prod_{j=1}^{m}\|\widetilde{S}_{j}f_{j}\|_{L^{p_{0j}}_{{w}_{0j}}(\Omega)}.$
Consequently, $\widetilde{T}\vec{f}$ is well-defined for all $\vec{f}\in\prod_{j=1}^{\infty}\widetilde{S}_{j}^{-1}(L^{\mathbf{p}_{j}}_{\boldsymbol{{w}}_{j}}(\Omega))$ by Theorem 3.4 with
$\|\widetilde{T}\vec{f}\|_{L^{\mathbf{q}}_{\boldsymbol{{v}}}(\Omega)}\leq\phi_{\vec{s}_{0},\vec{r}_{0},\vec{\mathbf{s}}}([\vec{\boldsymbol{{w}}}]_{(\vec{\mathbf{s}},\vec{\mathbf{r}})})\prod_{j=1}^{m}\|\widetilde{S}_{j}f_{j}\|_{L^{\mathbf{p}_{j}}_{\boldsymbol{{w}}_{j}}(\Omega)},$
@@
    $\|{\mathcal{M}}_{{\mathcal{U}}_{\Lambda_{i}}}\|_{L^{t_{i}}_{{w}_{t_{i}}}(\Omega_{\Lambda_{i}})\to L^{t_{i}}_{{v}_{t_{i}}}(\Omega_{\Lambda_{i}})}\leq\psi_{i}([{w}_{i},{v}_{i}]_{(s_{i},r_{i})}),$
@@
    There is $\mathcal{T}\subseteq\{T:\bigcup_{({w}_{0},{v}_{0})\in W}L^{p_{0}}_{{w}_{0}}(\Omega)\to L^{0}(\Omega)\text{ is linear}\}$ and a function $\phi:\mathbb{R}^{m}\to\mathbb{R}$ increasing in each component such that for all $({w}_{0},{v}_{0})\in W$, all $T\in{\mathcal{T}}$ and all $f\in L^{p_{0}}_{{w}_{0}}(\Omega)$ it holds
    $\|Tf\|_{L^{q_{0}}_{{v}_{0}}(\Omega)}\leq\phi([{w}_{0},{v}_{0}]_{s_{0},r_{0})})\|f\|_{L^{p_{0}}_{{w}_{0}}(\Omega)}.$
Then there is a function $\phi_{s_{0},r_{0},\boldsymbol{\gamma}}:\mathbb{R}^{\ell}\to\mathbb{R}$ increasing in each component such that $Tf$ is well-defined for all $T\in{\mathcal{T}}$ and $f\in L^{\mathbf{p}}_{\boldsymbol{{w}}}(\Omega)$, and ${\mathcal{T}}\subseteq{\mathcal{L}}(L^{\mathbf{p}}_{\boldsymbol{{w}}}(\Omega),L^{\mathbf{q}}_{\boldsymbol{{v}}}(\Omega))$ is $\mathcal{R}$}-bounded with
@@
By Theorem 3.6 applied with $m=1$, we have
$\|(T_{k}f_{k})_{k\in\mathbb{N}}\|_{L^{\mathbf{q}}_{\boldsymbol{{v}}}(\Omega;\ell^{2})}\leq\phi_{\vec{s_{0}},\vec{r_{0}},\vec{\boldsymbol{\gamma}}}([\vec{\boldsymbol{{w}}},\vec{\boldsymbol{{v}}}]_{\vec{\mathbf{s}},\vec{\mathbf{r}}})\|(f_{k})_{k\in\mathbb{N}}\|_{L^{\mathbf{p}}_{\boldsymbol{{w}}}(\Omega;\ell^{2})}$
for all $(T_{k})_{k\in\mathbb{N}}\subseteq\mathcal{T}$ and $(f_{k})_{k\in\mathbb{N}}\subseteq L^{\mathbf{p}}_{\boldsymbol{{w}}}(\Omega;\ell^{u_{2}})$. By our assumptions on $\mathbf{q}$ and $\mathbf{p}$, both $L^{\mathbf{p}}_{\boldsymbol{{w}}}(\Omega)$ and $L^{\mathbf{q}}_{\boldsymbol{{v}}}(\Omega)$ are Banach spaces. Since $L^{\mathbf{q}}_{\boldsymbol{{v}}}(\Omega)=L^{\mathbf{q}}(\Omega,\otimes{v}_{i}^{q_{i}}\,\mathrm{d}\mu)$ has finite cotype with cotype-bound independent of the weight by [9, Corollary 7.1.5], the $\mathcal{R}$}-bound follows by [9, Theorem 8.1.3(iii)]. ∎

### v2

@@ 3 Limited-Range Multilinear Off-Diagonal Mixed-Norm Extrapolation
$\|\widetilde{T}\vec{f}\|_{L^{\tilde{\mathbf{q}}}_{\tilde{\boldsymbol{{w}}}}(\Omega_{\Lambda_{2}})}\leq\kappa^{(\ell-1)/\alpha}\tilde{\phi}((\tilde{\phi}_{j}(\kappa^{\prime},[\tilde{\boldsymbol{{w}}}_{j},\tilde{\boldsymbol{{v}}}_{j}]_{\tilde{\mathbf{s}}_{j},\tilde{\mathbf{r}}_{j}}))_{j\in J})\prod_{j\in J}\|\widetilde{S}_{j}f_{j}\|_{L^{\tilde{\mathbf{p}}_{j}}_{\tilde{\boldsymbol{{v}}}_{j}}(\Omega_{\Lambda_{2}})},$ (20)
@@
    The proof of Theorem 3.4 allows to track the exact definition of $\phi_{\vec{s_{0}},\vec{r_{0}},\vec{\boldsymbol{\gamma}}}$ recursively. In particular, assume that there is $c:(1,\infty]\to[1,\infty)$ such that for all $p\in(1,\infty]$, $i\in I$, and weights w{w} on $\Omega_{\Lambda_{i}}$ with ${w}]_{p}^{{\mathcal{U}}_{\Lambda_{i}}}<\infty$, we have that ${\mathcal{M}}_{{\mathcal{U}}_{\Lambda_{i}}}$ is bounded in $L^{p}_{{w}}(\Omega_{\Lambda_{i}})$ with
@@
    Then if $\vec{\mathbf{s}},\vec{\mathbf{r}}\in(0,\infty)^{\ell\times m}$, $\vec{\boldsymbol{{w}}}=\vec{\boldsymbol{{v}}}$, and $\kappa\in(1,\infty)$, an iteration of the bound (11) in Remark 2.7 shows that we may choose $\phi_{\vec{s_{0}},\vec{r_{0}},\vec{\boldsymbol{\gamma}}}([\vec{\boldsymbol{{w}}},\vec{\boldsymbol{{v}}}]_{\vec{\mathbf{s}},\vec{\mathbf{r}}}):=\kappa^{\ell/\alpha}\phi(\vec{C})$ with
@@
    Theorem 3.4 also contains a weak-type extrapolation result by an argument of Grafakos and Martell [9]: One may exchange $\|T\vec{f}\|_{L^{q_{0}}_{{v}_{0}}(\Omega)}$ by $\|T\vec{f}\|_{L^{q_{0},\infty}_{{v}_{0}}(\Omega)}$ in (16) if one also exchanges $\|T\vec{f}\|_{L^{\mathbf{q}}_{\boldsymbol{{v}}}(\Omega)}$ by $\|T\vec{f}\|_{L^{\mathbf{q},\infty}_{\boldsymbol{{v}}}(\Omega)}$ in (17). Here, the weak mixed-norm space $L^{\mathbf{p},\infty}(\Omega)$ is given as all $f\in L^{0}(\Omega)$ such that
@@
    where $E_{\lambda}:=\{x\in\Omega\ \lvert\ |f|>\lambda\}$, and $f\in L^{\mathbf{p},\infty}_{\boldsymbol{{v}}}(\Omega)$ if $f{v}\in L^{\mathbf{p},\infty}(\Omega)$, where v{v} is the tensor weight induced by $\boldsymbol{{v}}$. To see the weak-type extrapolation, consider for $\lambda>0$ the set $E_{\lambda}:=\{x\in\Omega\ \lvert\ |T\vec{f}(x)|>\lambda\}$ and define $T_{\lambda}\vec{f}:=\lambda\mathds{1}_{E_{\lambda}}$. Applying Theorem 3.4 with $T\mapsto T_{\lambda}$, the assertion is obtained in the same way as in [9, Theorem 6.1].
@@
$\mathcal{T}\subseteq\{T:\prod_{j\in J}\bigcup_{({w}_{0j},{v}_{0j})\in W_{j}}S_{j}^{-1}(L^{p_{0j}}_{{v}_{0j}}(\Omega))\to L^{0}(\Omega)\}.$
If there is a function $\phi:\mathbb{R}^{m}\to\mathbb{R}$ increasing in each component such that for all $T\in{\mathcal{T}}$, $(\vec{{w}}_{0},\vec{{v}}_{0})\in\vec{W}:=W_{1}\times\cdots\times W_{m}$, and $\vec{f}\in\prod_{j\in J}S_{j}^{-1}(L^{\vec{p}_{0}}_{\vec{{v}}_{0}}(\Omega))$ it holds
$\|T\vec{f}\|_{L^{q_{0}}_{{w}_{0}}(\Omega)}\leq\phi([\vec{{w}}_{0},\vec{{v}}_{0}]_{(\vec{s}_{0},\vec{r}_{0})})\prod_{j\in J}\|S_{j}f_{j}\|_{L^{p_{0j}}_{{v}_{0j}}(\Omega)},$ (21)
then for all $T\in\mathcal{T}$ and all $\vec{f}\in\prod_{j\in J}S_{j}^{-1}(L^{\mathbf{p}_{j}}_{\boldsymbol{{v}}_{j}}(\Omega))$, $T\vec{f}$ is well defined, and
$\|(T_{k}(f_{1k},\ldots,f_{jk}))_{k\in\mathbb{N}}\|_{L^{\mathbf{q}}_{\boldsymbol{{w}}}(\Omega;\ell^{u_{0}})}\leq\phi_{\vec{s_{0}},\vec{r_{0}},\vec{\boldsymbol{\gamma}}}([\vec{\boldsymbol{{w}}},\vec{\boldsymbol{{v}}}]_{\vec{\mathbf{s}},\vec{\mathbf{r}}})\prod_{j\in J}\|(f_{jk})_{k\in\mathbb{N}}\|_{L^{\mathbf{p}_{j}}_{\boldsymbol{{v}}_{j}}(\Omega;\ell^{u_{0j}})},$
for all $(T_{k})_{k\in\mathbb{N}}\subseteq\mathcal{T}$ and $(f_{jk})_{k\in\mathbb{N}}\subseteq V_{j}$ with $(S_{j}f_{jk})_{k\in\mathbb{N}}\in L^{\mathbf{p}_{j}}_{\boldsymbol{{v}}_{j}}(\Omega;\ell^{u_{0j}})$ for all $j\in J$. The function $\phi_{\vec{s_{0}},\vec{r_{0}},\vec{\boldsymbol{\gamma}}}$ can be taken identical to the one in Theorem 3.4.
@@
$\widetilde{V}_{j}:=\{(f_{jk})_{k\in\mathbb{N}}\subseteq V_{j}\ \lvert\ (S_{j}f_{jk})_{k\in\mathbb{N}}\in\bigcup_{({w}_{0j},{v}_{0j})\in W_{j}}L^{p_{0j}}_{{v}_{0j}}(\Omega;\ell^{u_{0j}})\}$
@@
We claim that $\widetilde{S}_{j}^{-1}(L^{\mathbf{p}_{j}}_{\boldsymbol{{v}}_{j}}(\Omega))=\{(f_{jk})_{k\in\mathbb{N}}\subseteq V_{j}\ \lvert\ (S_{j}f_{jk})_{k\in\mathbb{N}}\in L^{\mathbf{p}_{j}}_{\boldsymbol{{v}}_{j}}(\Omega;\ell^{u_{0j}})\}$ for all $j\in J$. The inclusion $\subseteq$” follows directly from the definition of $\widetilde{S}_{j}$. For $\supseteq$” let $(f_{jk})_{k\in\mathbb{N}}$ be in the set on the right-hand side and consider the function $h:=\|(S_{j}f_{jk})_{k\in\mathbb{N}}\|_{\ell^{u_{0j}}}\in L^{\mathbf{p}_{j}}_{\boldsymbol{{v}}_{j}}(\Omega)$. Then $h\in\bigcup_{({w}_{0j},{v}_{0j})\in W_{j}}L^{p_{0j}}_{{v}_{0j}}(\Omega)$ by Lemma 3.3, that is $(f_{jk})_{k\in\mathbb{N}}\subseteq\widetilde{V}_{j}$. Hence we may apply $\widetilde{S}_{j}$ to $(f_{jk})_{k\in\mathbb{N}}$ to see that $\widetilde{S}_{j}(f_{jk})_{k\in\mathbb{N}}=h\in L^{\mathbf{p}_{j}}_{\boldsymbol{{v}}_{j}}(\Omega)$, i.e. $(f_{jk})_{k\in\mathbb{N}}\in\widetilde{S}_{j}^{-1}(L^{\mathbf{p}_{j}}_{\boldsymbol{{v}}_{j}}(\Omega))$.
@@
$\|(S_{j}f_{jk})_{k}\|_{\ell^{u_{0j}}(L^{p_{0j}}_{{v}_{0j}}(\Omega))}\leq\|\widetilde{S}_{j}f_{j}\|_{L^{p_{0j}}_{{v}_{0j}}(\Omega)}.$
@@
$\|\widetilde{T}\vec{f}$ $\|_{L^{q_{0}}_{\vec{{w}}_{0}}(\Omega)}\leq\|(T_{k}\vec{f}_{k})_{k}\|_{\ell^{u_{0}}(L^{q_{0}}_{\vec{{w}}_{0}}(\Omega))}\leq\phi([{\vec{{w}}_{0}}]_{(\vec{s}_{0},\vec{r}_{0})})\left\|\prod_{j=1}^{m}\|(S_{j}f_{jk})_{k}\|_{L^{p_{0j}}_{{v}_{0j}}(\Omega)}\right\|_{\ell^{u_{0}}}$
$\leq\phi([{\vec{{w}}_{0}}]_{(\vec{s}_{0},\vec{r}_{0})})\prod_{j=1}^{m}\|(S_{j}f_{jk})_{k}\|_{\ell^{u_{0j}}(L^{p_{0j}}_{{v}_{0j}}(\Omega))}\leq\phi([{\vec{{w}}_{0}}]_{(\vec{s}_{0},\vec{r}_{0})})\prod_{j=1}^{m}\|\widetilde{S}_{j}f_{j}\|_{L^{p_{0j}}_{{v}_{0j}}(\Omega)}.$
Consequently, $\widetilde{T}\vec{f}$ is well-defined for all $\vec{f}\in\prod_{j=1}^{\infty}\widetilde{S}_{j}^{-1}(L^{\mathbf{p}_{j}}_{\boldsymbol{{v}}_{j}}(\Omega))$ by Theorem 3.4 with
$\|\widetilde{T}\vec{f}\|_{L^{\mathbf{q}}_{\boldsymbol{{w}}}(\Omega)}\leq\phi_{\vec{s}_{0},\vec{r}_{0},\vec{\mathbf{s}}}([\vec{\boldsymbol{{w}}}]_{(\vec{\mathbf{s}},\vec{\mathbf{r}})})\prod_{j=1}^{m}\|\widetilde{S}_{j}f_{j}\|_{L^{\mathbf{p}_{j}}_{\boldsymbol{{v}}_{j}}(\Omega)},$
@@
    $\|{\mathcal{M}}_{{\mathcal{U}}_{\Lambda_{i}}}\|_{L^{t_{i}}_{{v}_{t_{i}}}(\Omega_{\Lambda_{i}})\to L^{t_{i}}_{{w}_{t_{i}}}(\Omega_{\Lambda_{i}})}\leq\psi_{i}([{w}_{i},{v}_{i}]_{(s_{i},r_{i})}),$
@@
    There is $\mathcal{T}\subseteq\{T:\bigcup_{({w}_{0},{v}_{0})\in W}L^{p_{0}}_{{w}_{0}}(\Omega)\to L^{0}(\Omega)\text{ is linear}\}$ and a function $\phi:\mathbb{R}^{m}\to\mathbb{R}$ increasing in each component such that for all $({w}_{0},{v}_{0})\in W$, all $T\in{\mathcal{T}}$ and all $f\in L^{p_{0}}_{{v}_{0}}(\Omega)$ it holds
    $\|Tf\|_{L^{q_{0}}_{{w}_{0}}(\Omega)}\leq\phi([{w}_{0},{v}_{0}]_{s_{0},r_{0})})\|f\|_{L^{p_{0}}_{{v}_{0}}(\Omega)}.$
Then there is a function $\phi_{s_{0},r_{0},\boldsymbol{\gamma}}:\mathbb{R}^{\ell}\to\mathbb{R}$ increasing in each component such that $Tf$ is well-defined for all $T\in{\mathcal{T}}$ and $f\in L^{\mathbf{p}}_{\boldsymbol{{v}}}(\Omega)$, and ${\mathcal{T}}\subseteq{\mathcal{L}}(L^{\mathbf{p}}_{\boldsymbol{{v}}}(\Omega),L^{\mathbf{q}}_{\boldsymbol{{w}}}(\Omega))$ is $\mathcal{R}$}-bounded with
@@
By Theorem 3.6 applied with $m=1$, $V:=L^{0}(\Omega)$ and $S:=\operatorname{\mathsf{id}}$, we have
$\|(T_{k}f_{k})_{k\in\mathbb{N}}\|_{L^{\mathbf{q}}_{\boldsymbol{{w}}}(\Omega;\ell^{2})}\leq\phi_{\vec{s_{0}},\vec{r_{0}},\vec{\boldsymbol{\gamma}}}([\vec{\boldsymbol{{w}}},\vec{\boldsymbol{{v}}}]_{\vec{\mathbf{s}},\vec{\mathbf{r}}})\|(f_{k})_{k\in\mathbb{N}}\|_{L^{\mathbf{p}}_{\boldsymbol{{v}}}(\Omega;\ell^{2})}$
for all $(T_{k})_{k\in\mathbb{N}}\subseteq\mathcal{T}$ and $(f_{k})_{k\in\mathbb{N}}\subseteq L^{\mathbf{p}}_{\boldsymbol{{w}}}(\Omega;\ell^{u_{2}})$. By our assumptions on $\mathbf{q}$ and $\mathbf{p}$, both $L^{\mathbf{p}}_{\boldsymbol{{v}}}(\Omega)$ and $L^{\mathbf{q}}_{\boldsymbol{{w}}}(\Omega)$ are Banach spaces. Since $L^{\mathbf{q}}_{\boldsymbol{{w}}}(\Omega)=L^{\mathbf{q}}(\Omega,\otimes{w}_{i}^{q_{i}}\,\mathrm{d}\mu)$ has finite cotype with cotype-bound independent of the weight by [10, Corollary 7.1.5], the $\mathcal{R}$}-bound follows by [10, Theorem 8.1.3(iii)]. ∎